Let V be the solution space of the following homogeneous linear system: x1 − x2 − 2x3 + 2x4 − 3x5 = 0 x1 − x2 − x3 + x4 − 2x5 = 0. (a) (2 points) Find a basis S of V and write down the dimension of V.
(b) (3 points) Finda subspace W of R5 suchthat W contains V anddim(W) = 4. Justify your answer.
1
Expert's answer
2020-03-17T12:46:55-0400

"x_1-x_2-2x_3+2x_4-3x_5=0"

"x_1-x_2-x_3+x_4-2x_5=0"

Solving both of them we get;

"x_1-x_2=2x_3-2x_4+3x_5"

"x_1-x_2=x_3-x_4+2x_5"

"2x_3-2x_4+3x_5=x_3-x_4+2x_5"

"x_4=x_3+x_5"

"x_1-x_2=x_3-x_4+2x_5"

"x_1-x_2=x_5"

"x_1=x_2+x_5"

"\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3\\
x_4\\
x_5
\end{pmatrix}=\begin{pmatrix}
 x_2+x_5 \\
 x_2 \\
 x_3\\
 x_3+x_5\\
 x_5
\end{pmatrix}"

"\begin{pmatrix}
 x_2+x_5 \\
 x_2 \\
 x_3\\
 x_3+x_5\\
 x_5
\end{pmatrix}=\begin{pmatrix}
 x_2\\
 x_2\\
0\\
0\\
0
\end{pmatrix}+\begin{pmatrix}
 x_5\\
 0\\
0\\
x_5\\
x_5
\end{pmatrix}+\begin{pmatrix}
 0\\
 0\\
x_3\\
x_3\\
0
\end{pmatrix}"

"\begin{pmatrix}
 x_2+x_5 \\
 x_2 \\
 x_3\\
 x_3+x_5\\
 x_5
\end{pmatrix}=\begin{pmatrix}
 1\\
 1\\
0\\
0\\
0
\end{pmatrix}x_2+\begin{pmatrix}
 1\\
 0\\
0\\
1\\
1
\end{pmatrix}x_5+\begin{pmatrix}
 0\\
 0\\
1\\
1\\
0
\end{pmatrix}x_3"

"{\begin{pmatrix}
 1\\
 1\\
0\\
0\\
0
\end{pmatrix},\begin{pmatrix}
 1\\
 0\\
0\\
1\\
1
\end{pmatrix},\begin{pmatrix}
 0\\
 0\\
1\\
1\\
0
\end{pmatrix}}" Is the basis of V.

Dimension of V is 3.

(b)If W is a solution space of homogenous system of equation "x_1-x_2-2x_3+2x_4-3x_5=0"

"x_1=x_2+2x_3-2x_4+3x_5"

"\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3\\
x_4\\
x_5
\end{pmatrix}=\begin{pmatrix}
 x_2+2x_3-2x_4+3x_5 \\
 x_2 \\
 x_3\\
 x_4\\
 x_5
\end{pmatrix}"

"\begin{pmatrix}
 x_2+2x_3-2x_4+3x_5 \\
 x_2 \\
 x_3\\
 x_4\\
 x_5
\end{pmatrix}=\begin{pmatrix}
 x_2\\
 x_2\\
0\\
0\\
0
\end{pmatrix}+\begin{pmatrix}
 2 x_3\\
 0\\
x_3\\
0\\
0
\end{pmatrix}+\begin{pmatrix}
 -2x_4\\
 0\\
0\\
x_4\\
0
\end{pmatrix}+\begin{pmatrix}
 3x_5\\
 0\\
0\\
0\\
x_5
\end{pmatrix}"

"\begin{pmatrix}
 x_2+2x_3-2x_4+3x_5 \\
 x_2 \\
 x_3\\
 x_4\\
 x_5
\end{pmatrix}=\begin{pmatrix}
 1\\
 1\\
0\\
0\\
0
\end{pmatrix}x_2+\begin{pmatrix}
 2 \\
 0\\
1\\
0\\
0
\end{pmatrix}x_3+\begin{pmatrix}
 -2\\
 0\\
0\\
1\\
0
\end{pmatrix}x_4+\begin{pmatrix}
 3\\
 0\\
0\\
0\\
1
\end{pmatrix}x_5"

"{\begin{pmatrix}
 1\\
 1\\
0\\
0\\
0
\end{pmatrix},\begin{pmatrix}
 2 \\
 0\\
1\\
0\\
0
\end{pmatrix} ,\begin{pmatrix}
- 2 \\
 0\\
0\\
1\\
0
\end{pmatrix},\begin{pmatrix}
 3 \\
 0\\
0\\
0\\
1
\end{pmatrix}}" Is the basis of W.

Dimension of vector space W is 4.